# Gap solitons of the Wannier and Bloch types in spin-orbit-coupled Bose-Einstein condensates with a moiré lattice

### v1

@@ Abstract
Gap solitons (GSs) bifurcating from flat bands, which may be represented in terms of Wannier functions, have garnered significant interest due to their strong localization with extremely small norms. Moiré lattices (MLs), with multiple flat bands, offer an appropriate platform for creating such solitons. We explore the formation mechanism and stability of GSs in spin-1 Bose-Einstein condensates under the combined action of the Rashba spin-orbit coupling (SOC) and an ML potential. We identify five Wannier-type GS families bifurcating from the lowest five energy bands in the spectrum induced by the ML with sufficiently large period and depth. These fundamental GSs serve as basic elements for constructing more complex Wannier-type GS states. Reducing the lattice period and depth triggers a transition from the Wannier-type GSs to ones of the Bloch type, the latter exhibiting higher norm thresholds and pronounced spatial broadening near edges of the energy bands. In addition to tuning the lattice-potential parameters, adjusting the SOC strength can also modulate the flatness of energy bands and enhance the localization of gap solitons, enabling reversible transitions between the GSs of the Wannier and Bloch types. Distinctive properties of GSs in the quasiperiodic ML are uncovered too. Thus, we propose the theoretical foundation for the creation of and manipulations with strongly localized GSs.
preprint: APS/123-QED
@@ I Introduction
Gap solitons (GSs) are nonlinear localized states populating spectral gaps induced by spatially periodic or quasiperiodic lattice potentials. GSs have been extensively studied in many fields, especially nonlinear optics and Bose-Einstein condensates (BECs) [1, 2, 3]. In particular, in two dimensions (2D) families of GS modes have been studied under the action of various potentials representing triangular [4, 5], square-shaped [6, 7, 8], hexagonal [9, 10], ring-shaped [11, 12] and quasiperiodic [13]. GSs are classified into two distinct types according to their shapes. A prevalent one is the Bloch type (BT) [14]. It features widely expanding GS wave functions, which asymptotically approach the corresponding linear Bloch states when the GS is created near edges of the corresponding energy bands, that are adjacent to the spectral gap hosting the GS. The BT GSs exist above a large norm threshold and are challenging to excite. GSs of the other type bifurcate from nearly-flat bands, and can be excited even with a very small norm. Being approximated by Wannier functions, rather than the Bloch ones, they are referred to as Wannier-type (WT) GSs [15, 16]. Owing to their tight localization and dynamical stability, the WT species of GSs offers applications to quantum information and topological quantum computing [17, 18].
@@
As said above, the emergence of WT GSs is intrinsically linked to the existence of flat bands in the respective linearized system. For conventional lattices, realizing such flat bands typically requires extremely deep potential wells, yielding only a limited number of effectively flat bands and constraining further advancement in this field. Recently, moiré lattices (MLs), built as a superposition of twos conventional 2D lattices with a special rotation (twist) angle between them, have drawn much interest, revealing intriguing phenomena, such as the unconventional superconductivity [19], fractal energy spectra [20], and the localization-delocalization transition [21, 22]. MLs exhibit a distinctive band-gap structure characterized by the alternation of multiple flat bands and wide gaps, providing a unique platform for exploring novel species of WT GSs. Initial studies of GSs in MLs were primarily conducted in terms of photonic systems, where the MLs are imprinting in photorefractive crystals by optical induction [23]. Various types of GSs have been thus identified, including multifrequency [24], multipole [25, 26], and vortex [27] solitons. With the realization of tunable twisted-bilayer optical lattices in BECs [28], the studies of ML-supported GSs has been extended to single-component [29] and two-component [30, 31] matter-wave systems. Most current studies focus on GSs bifurcating from the lowest energy band, while systematic investigations are lacking for GS modes that may bifurcate from higher energy bands.
@@
The spin-orbit coupling (SOC) introduces an additional tool for the work with matter-wave systems. Mathematically represented by the first-order differential operator, SOC shows the promise for controlling flat bands [32, 33, 34]. In other contexts, SOC suppresses the collapse of 2D [35, 36, 37, 38] and 3D [39] solitons, also enabling the creation of diverse types of “exotic” solitons [40, 41, 42, 43, 44, 45]. However, specific forms and properties of GSs were not addressed under the combined effect of SOC and MLs, especially in three-component BEC systems, with their large set of tunable parameters.
@@
In this work, we focus on a effectively 2D matter-wave system with SOC of the Rashba type [46, 47] applied along with an ML. For periodic MLs, we systematically investigate the existence and stability of GSs bifurcating from the lowest five energy bands for different interaction strengths and lattice periods. The interaction determines whether the GS bifurcates to the gap above or below the corresponding flat band, while the lattice period determines the type of the emerging GS. For the large lattice period, the lowest five energy bands are effectively flat, and we find five fundamental WT GS families bifurcating from these flat bands. The coherent superposition of the fundamental solitons enables the creation of more complex WT-GS modes. Conversely, for the small period, BT GSs emerge, with phase profiles congruent to those of WT GSs. We find that SOC plays a crucial role in flattening the energy bands, thereby facilitating the transition of WT GSs to BT GSs. Finally, we extend the analysis to quasiperiodic MLs.
@@
The presentation is structured as follows. In Sec. II we describe the mean-field model for the spin-1 BEC with the Rashba SOC loaded into the ML. In Sec. III we report systematically produced analytical and numerical results for periodic and quasiperiodic MLs, including the linear bandgap spectrum, two (WT and BT) GSs species, and effects of various parameters. The paper is concluded by Sec. IV.
@@ II Model
We consider a quasi-2D spin-1 BEC, with coordinates $\left(x,y\right)$, under the action of the Rashba SOC and ML. In the framework of the mean-field approximation, the system is governed by the three-component Gross-Pitaevskii equation [48, 49], which is written here in the scaled form,
@@
where the units of the energy, time, and length are $\hbar\omega_{z}$, $\omega_{z}^{-1}$, and $l_{z}=\sqrt{\hbar/(m\omega_{z})}$, respectively. $m$ is the atomic mass, and $\omega_{z}$ is the trapping frequency applied to the BEC layer in the transverse direction. $\psi_{j}$ and $\rho_{j}=\left|\psi_{j}\right|^{2}$ (with $j=\pm 1,0$) are three-component wave functions and densities, respectively. $\rho=\sum_{j}\left|\psi_{j}\right|^{2}$ is the total density. Constants of the mean-field and spin-exchange interactions are $c_{0}=2\sqrt{2\pi}(a_{0}+2a_{2})/(3l_{z})$ and $c_{2}=2\sqrt{2\pi}(a_{2}-a_{0})/(3l_{z})$, respectively, where $a_{0}$ and $a_{2}$ are two-body s-wave scattering lengths for the total spin 0 and $2$ [50]. Further, $\gamma$ is the strength of Rashba SOC, which is represented by operators $\partial_{\pm}=i\partial_{y}\pm\partial_{x}$. The ML is represented by its potential, constructed by overlapping two square-shaped optical lattices with a twist angle $\theta$, between them,
@@
In the experiment, the Rashba SOC can be induced by the Raman coupling in the square optical lattice [51]. The ML system is produced by adding another square optical lattice (which is rotated by angle $\theta$). Parameters ($\gamma$, $a$, $V_{0}$) may be controlled by adjusting laser beams illuminating the condensate. Furthermore, the nonlinear parameters $c_{0}$ and $c_{2}$ can be adjusted by means of the Feshbach-resonance technique [52, 53].
@@ III Results and discussion
III.1 The bandgap structures
When the twisted angle $\theta$ takes a Pythagorean value, the two square sublattices form moiré patterns with the periodic translational symmetry, the period being $\sqrt{\bar{c}}a$. We focus on two Pythagorean angles, $\theta=\arctan(3/4)$ and $\theta=\arctan(5/12)$, with the profiles of the corresponding MLs with $V_{0}=4$ are shown in Figs. 1(a) and 1(b), which are affected by the lattice period $a$.
Figure 1: The ML profiles for twisted angles (a) $\theta=\arctan(3/4)\approx 36.9^{\circ}$ and (b) $\theta=\arctan(5/12)\approx 22.6^{\circ}$. The black arrows represent the lattice vectors. The bandgap structures along the high-symmetry lines for the MLs with (c) $\theta=\arctan(3/4)$ and (d) $\theta=\arctan(5/12)$ when $a=0.5\pi$ and $\gamma=0.5$. The box and red arrows in (c) represent the first Brillouin zone and the high-symmetry lines. (e)-(f) The effect of $\gamma$ and $a$ on the ML bandgap structure with $\theta=\arctan(3/4)$. These blue (white) areas represent energy bands (gaps). Here, $V_{0}=4$.
The perpendicular lattice vectors (black arrows) imply that the reciprocal lattice vectors are also perpendicular to each other. The length of the reciprocal lattice vector is $2\pi/(\sqrt{\bar{c}}a)$, which is also the side width of the first Brillouin zone, as shown in the subplot of Fig. 1(c). The ML periodicity makes it possible to calculate the linear bandgap spectrum of the system by means of the linear Bloch theory, which is the necessary step in the study of GSs.
@@ III.1 The bandgap structures
where $\mu_{n}$ is the chemical potential of the energy band with index $n$. $k_{x}$ and $k_{y}$ are the Bloch quasi-wavenumbers. By substituting Eq. (4) in the linearized equation (1), we derive the equations that the eigenfunctions $\phi^{(n)}_{j}$ satisfy
@@
Solving Eq. (5) by mean of the Fourier collocation method, we can obtain the chemical potential as the function of the quasi-wavenumbers, $\mu_{n}(k_{x},k_{y})$, and the corresponding eigenfunctions $\phi_{j}^{(n)}(x,y,k_{x},k_{y})$ in the first Brillouin zone. Figures 1(c) and 1(d) show the band-gap structures along the high-symmetry lines for the MLs with $\theta=\arctan(3/4)$ and $\theta=\arctan(5/12)$, respectively. Comparing the lowest three energy bands in these two cases, we find that the energy bands for $\theta=\arctan(5/12)$ are relatively flat and have lower values of $\mu_{n}$. However, the energy gaps for $\theta=\arctan(3/4)$ are much wider than those for $\theta=\arctan(5/12)$, especially for the second finite energy gap. The presence of multiple and wide energy gaps enables the system to support various GS types.
The depth and period of the lattice potential significantly affect the bandgap structure, a deeper lattice potential resulting in narrower bands and wider band gaps. We consider the effect of lattice period $a$ on the bandgap structure for $\theta=\arctan(5/12)$, $V_{0}=4$, and $\gamma=0.5$, as shown in Fig. 1(e). As the lattice period $a$ increases, the chemical potential $\mu$ of each energy band decreases. Simultaneously, the increase in $a$ leads, first, to merger of the high-energy bands, followed by splitting into narrower (nearly flat) bands, accompanied by the disappearance and reemergence of the gaps which separate the bands. The emergence of multiple flat bands here is an ML-produced effect.
We also adddress the effect of the SOC strength $\gamma$ on the bandgap structure fixing given $\theta=\arctan(5/12)$, $V_{0}=4$, and $a=0.5$, as shown in Fig. 1(f). Similar to the effect of $a$, the increase in $\gamma$ leads to a decrease in the chemical potential of each energy band, but at a faster rate. Due to the varying rates of the decrease for each energy band, the widths of the energy gaps are affected, and an additional energy gap forms.
@@ III.2 Wannier-type and Bloch-type gap solitons
We have found bright solitons, i.e., GSs populating the gaps. If both interaction coefficients $c_{0}$ and $c_{2}$ are negative (which corresponds to the attractive nonlinearity), the 2D system suffers the collapses, as might expected [54, 55], which inhibits the formation of GSs. On the other hand, if both $c_{0}$ and $c_{2}$ are positive (which corresponds to the repulsive nonlinearity), the broadening of wave functions induced by dispersion cannot be balanced by the repulsive interactions. GSs can only exist in a large-scale potential well (binding all atoms) of the ML, which is similar to the harmonic-oscillator trapping potential [56]. Here, we produce GS solutions of Eq. (1) in the most interesting case of $c_{0}c_{2}<0$, i.e., two cases $c_{0}<0$, $c_{2}>0$ (antiferromagnetic) and $c_{0}>0$, $c_{2}<0$ (ferromagnetic) [48]. Here, we fix $\theta=\arctan(3/4)$ to deal with broader energy gaps.
GS solutions $\varphi_{j}(x,y)$ satisfy the stationary-state equations, which are obtained by substituting $\psi_{j}(x,y,t)=\varphi_{j}(x,y)e^{i\mu t}$ into Eq. (1)
@@
GS solutions to Eq. (6) were obtained using the numerical squared-operator iteration method [57], with the computational grid $(x,y)\in[-12\pi,+12\pi]$ and the number of discrete points is $1024\times 1024$. For the ML with $\theta=\arctan(3/4)$, $a=2\pi$, and $V_{0}=4$, the low-energy bands exhibit flat dispersion, and several gaps are present. We have found five types of GSs, setting $\gamma=0.5$, $c_{0}=-1$, and $c_{2}=1.5$ (the antiferromagnetic system). The density profiles of the five GS types are plotted in Figs. 2(a)-2(e), featuring ring-shaped and multi-hump patterns.
Figure 2: WT (Wannier-type) GS families bifurcating from the lowest five flat bands for the ML with $\theta=\arctan(3/4)$, $a=2\pi$, and $V_{0}=4$. Other parameters are $\gamma=0.5$, $c_{0}=-1$ and $c_{2}=1.5$ (the antiferromagnetic system). (a)-(e) The density profiles of three components $\varphi_{\pm 1}$ (the first row) and $\varphi_{0}$ (the second row) for five GS families. The corresponding phases are shown in the corresponding subfigures, where the phase of $\varphi_{+1}$ is in the left subfigure. The drawing grid is $(x,y)\in[-3,3]$. (f) Norm NN versus the chemical potential $\mu$ for the five GS families. The solid (dashed) parts of these curves indicate that the corresponding GSs are stable (unstable). The black lines represent five flat bands with $\mu_{1}=1.6998$, $\mu_{2}=1.8065$, $\mu_{3}=3.5533$, $\mu_{4}=3.5808$, and $\mu_{5}=3.7703$. Here, the inset shows the norm curve of the GS family bifurcating from $\mu_{1}$, plotted near the band edge.
Note that the density profiles of the GS components $\psi_{+1}$ and $\psi_{-1}$ are identical, but they have different phases. In Fig. 2(f), we also present the relation between the soliton’s norm, $N=\int_{-\infty}^{+\infty}\sum_{j}|\varphi_{j}|^{2}\mathrm{d}x\mathrm{d}y$, and the chemical potential $\mu$ for the five GS families. The norm curves bifurcate at the chemical potential $\mu_{n}$ ($n=1,2,3,4,5$) of the corresponding flat bands and pass through the lower-energy flat bands towards $\mu\rightarrow$-$\infty$. In other words, the chemical potential of the GS families is
@@
where $\Delta\mu<0$, provided that $c_{0}<0$. It is easy to understood that the offset $\Delta\mu$ is determined by $c_{0}$ because these GS solutions are invariant with respect to the spin-flip symmetry (7), which implies that the nonlinear terms associated with $c_{2}$ are eliminated when substituting the GS solutions into Eq. (6). As $\mu$ increases, the density profiles of these five GS types remain almost unchanged (slightly widened); however, the norm demonstrates the trend of vanishing, $N\rightarrow 0$, at $\mu\rightarrow\mu_{n}$. In the actually simulated system, the GS norm does not strictly vanish, exceeding a very small threshold $\sim$ $10^{-8}$, as shown in the inset of Fig. 2. In addition, the square minimal unit cell of the ML gives rise to three other flat bands that are very close to those ($\mu_{2},\mu_{3},\mu_{4}$), and there exist three additional GS types (not shown here) whose density profiles are very similar to ones plotted in Figs. 2(b)-2(d), rotated by $\pi/4$. The norm curves of the two degenerate GS types are strongly overlapping. In fact, all these GSs are similar to Wannier functions $w_{j}^{(n)}(x,y)$ (in the central unit cell) corresponding to $\mu_{n}$, especially in the limit of small amplitudes, where the nonlinear terms may be treated as small perturbations, i.e.,
@@
where $S=\pi^{2}/(5a^{2})$ is the area of the first Brillouin zone [58]. The Wannier functions corresponding to the flat bands are eigenfunctions of the linear part of Eq. (1), with the eigenvalue of $\mu_{n}$. This is why these GS families bifurcate at the chemical potential $\mu_{n}$ corresponding to the flat bands. The GSs, which remain localized even at small values of the norm, can be well approximated by Wannier functions, therefore they are termed WT GSs.

### v2

@@ Abstract
Gap solitons (GSs) bifurcating from flat bands, which may be represented in terms of Wannier functions, have garnered significant interest due to their strong localization with extremely small norms. Moiré lattices (MLs), with multiple flat bands, offer an appropriate platform for creating such solitons. We explore the formation mechanism and stability of GSs in spin-1 Bose-Einstein condensates under the combined action of the Rashba spin-orbit coupling (SOC) and an ML potential. We identify five Wannier-type GS families bifurcating from the lowest five energy bands in the spectrum induced by the ML with sufficiently large period and depth. These fundamental GSs serve as basic elements for constructing more complex Wannier-type GS states. Reducing the lattice period and depth triggers a transition from the Wannier-type GSs to those of the Bloch type, the latter exhibiting higher norm thresholds and pronounced spatial broadening near edges of the energy bands. In addition to tuning the lattice-potential parameters, adjusting the SOC strength can also modulate the flatness of energy bands and enhance the localization of gap solitons, enabling reversible transitions between the GSs of the Wannier and Bloch types. Distinctive properties of GSs in the quasiperiodic ML are uncovered too. Thus, we propose the theoretical foundation for the creation of and manipulations with strongly localized GSs.
@@ I Introduction
Gap solitons (GSs) are nonlinear localized states populating spectral gaps induced by spatially periodic or quasiperiodic lattice potentials. GSs have been extensively studied in many fields, especially nonlinear optics and Bose-Einstein condensates (BECs) [1, 2, 3]. In particular, in two dimensions (2D) families of GS modes have been studied under the action of various potentials representing triangular [4, 5], square-shaped [6, 7, 8], hexagonal [9, 10], ring-shaped [11, 12], and quasiperiodic [13]. GSs are classified into two distinct types according to their shapes. A prevalent one is the Bloch type (BT) [14]. It features widely expanding GS wave functions, which asymptotically approach the corresponding linear Bloch states when the GS is created near edges of the corresponding energy bands that are adjacent to the spectral gap hosting the GS. The BT GSs exist above a large norm threshold and are challenging to excite. GSs of the other type bifurcate from nearly flat bands and can be excited even with a very small norm. Being approximated by Wannier functions rather than the Bloch ones, they are referred to as Wannier-type (WT) GSs [15, 16]. Owing to their tight localization and dynamic stability, the WT species of GSs offers applications to quantum information and topological quantum computing [17, 18].
@@
As said above, the emergence of WT GSs is intrinsically linked to the existence of flat bands in the respective linearized system. For conventional lattices, realizing such flat bands typically requires extremely deep potential wells, yielding only a limited number of effectively flat bands and constraining further advancement in this field. Recently, moiré lattices (MLs), built as a superposition of two conventional 2D lattices with a special rotation (twist) angle between them, have drawn much interest, revealing intriguing phenomena, such as unconventional superconductivity [19], fractal energy spectra [20], and the localization-delocalization transition [21, 22]. MLs exhibit a distinctive band-gap structure characterized by the alternation of multiple flat bands and wide gaps, providing a unique platform for exploring novel species of WT GSs. Initial studies of GSs in MLs were primarily conducted in terms of photonic systems, where the MLs are imprinting in photorefractive crystals by optical induction [23]. Various types of GSs have been thus identified, including multifrequency [24], multipole [25, 26], and vortex [27] solitons. With the realization of tunable twisted-bilayer optical lattices in BECs [28], the studies of ML-supported GSs have been extended to single-component [29] and two-component [30, 31] matter-wave systems. Most current studies focus on GSs bifurcating from the lowest-energy band, while systematic investigations are lacking for GS modes that may bifurcate from higher energy bands.
@@
The spin-orbit coupling (SOC) introduces an additional tool for the work with matter-wave systems. Mathematically represented by the first-order differential operator, SOC shows the promise for controlling flat bands [32, 33, 34]. In other contexts, SOC suppresses the collapse of 2D [35, 36, 37, 38] and 3D [39] solitons, also enabling the creation of diverse types of “exotic” solitons [40, 41, 42, 43, 44, 45]. However, specific forms and properties of GSs were not addressed under the combined effect of SOC and MLs, especially in three-component BEC systems with many tunable parameters.
@@
In this work we focus on a effectively 2D matter-wave system with SOC of the Rashba type [46, 47] applied along with an ML. For periodic MLs, we systematically investigate the existence and stability of GSs bifurcating from the lowest five energy bands for different interaction strengths and lattice periods. The interaction determines whether the GS bifurcates to the gap above or below the corresponding flat band, while the lattice period determines the type of the emerging GS. For the large lattice period, the lowest five energy bands are effectively flat, and we find five fundamental WT GS families bifurcating from these flat bands. The coherent superposition of the fundamental solitons enables the creation of more complex WT-GS modes. Conversely, for the small period, BT GSs emerge, with phase profiles congruent to those of WT GSs. We find that SOC plays a crucial role in flattening the energy bands, thereby facilitating the transition of WT GSs to BT GSs. Finally, we extend the analysis to quasiperiodic MLs.
@@
The presentation is structured as follows. In Sec. II we describe the mean-field model for the spin-1 BEC with the Rashba SOC loaded into the ML. In Sec. III we report systematically produced analytical and numerical results for periodic and quasiperiodic MLs, including the linear bandgap spectrum, two (WT and BT) GSs species, and effects of various parameters. The paper is concluded in Sec. IV.
@@ II Model
We consider a quasi-2D spin-1 BEC, with coordinates $\left(x,y\right)$, under the action of the Rashba SOC and ML. In the framework of the mean-field approximation, the system is governed by the three-component Gross-Pitaevskii equation [48, 49], which is written here in the scaled form:
@@
where the units of the energy, time, and length are $\hbar\omega_{z}$, $\omega_{z}^{-1}$, and $l_{z}=\sqrt{\hbar/(m\omega_{z})}$, respectively. $m$ is the atomic mass, and $\omega_{z}$ is the trapping frequency applied to the BEC layer in the transverse direction. $\psi_{j}$ and $\rho_{j}=\left|\psi_{j}\right|^{2}$ (with $j=\pm 1,0$) are three-component wave functions and densities, respectively. $\rho=\sum_{j}\left|\psi_{j}\right|^{2}$ is the total density with the norm $N$. Constants of the mean-field and spin-exchange interactions are $c_{0}=2\sqrt{2\pi}\mathcal{N}(a_{0}+2a_{2})/(3Nl_{z})$ and $c_{2}=2\sqrt{2\pi}\mathcal{N}(a_{2}-a_{0})/(3Nl_{z})$, respectively, where $a_{0}$ and $a_{2}$ are two-body s-wave scattering lengths for the total spin 0 and $2$ [50]. $\mathcal{N}$ is the number of atoms. Further, $\gamma$ is the strength of Rashba SOC, which is represented by operators $\partial_{\pm}=i\partial_{y}\pm\partial_{x}$. The ML is represented by its potential, constructed by overlapping two square-shaped optical lattices with a twist angle $\theta$ between them:
@@
In the experiment the Rashba SOC can be induced by the Raman coupling in the square optical lattice [51]. The ML system is produced by adding another square optical lattice (which is rotated by angle $\theta$). Parameters ($\gamma$, $a$, $V_{0}$) may be controlled by adjusting laser beams illuminating the condensate. Furthermore, the nonlinear parameters $c_{0}$ and $c_{2}$ can be adjusted by means of the Feshbach-resonance technique [52, 53].
@@ III Results and discussion
III.1 The band-gap structures
When the twisted angle $\theta$ takes a Pythagorean value, the two square sublattices form moiré patterns with the periodic translational symmetry, the period being $\sqrt{\bar{c}}a$. We focus on two Pythagorean angles, $\theta=\arctan(3/4)$ and $\theta=\arctan(5/12)$, with the profiles of the corresponding MLs with $V_{0}=4$ shown in Figs. 1(a) and 1(b), which are affected by the lattice period $a$.
Figure 1: The ML profiles for twisted angles (a) $\theta=\arctan(3/4)\approx 36.9^{\circ}$ and (b) $\theta=\arctan(5/12)\approx 22.6^{\circ}$. The black arrows represent the lattice vectors. The band-gap structures along the high-symmetry lines for the MLs with (c) $\theta=\arctan(3/4)$ and (d) $\theta=\arctan(5/12)$ when $a=0.5\pi$ and $\gamma=0.5$. The box and red arrows in (c) represent the first Brillouin zone and the high-symmetry lines. (e,f) The effect of $\gamma$ and $a$ on the ML band-gap structure with $\theta=\arctan(3/4)$. These blue (white) areas represent energy bands (gaps). Here, $V_{0}=4$.
The perpendicular lattice vectors (black arrows) imply that the reciprocal lattice vectors are also perpendicular to each other. The length of the reciprocal lattice vector is $2\pi/(\sqrt{\bar{c}}a)$, which is also the side width of the first Brillouin zone, as shown in the subplot of Fig. 1(c). The ML periodicity makes it possible to calculate the linear band-gap spectrum of the system by means of the linear Bloch theory, which is the necessary step in the study of GSs.
@@ III.1 The band-gap structures
where $\mu_{n}$ is the chemical potential of the energy band with index $n$. $k_{x}$ and $k_{y}$ are the Bloch quasi wave numbers. By substituting Eq. (4) in the linearized equation (1), we derive the equations that the eigenfunctions $\phi^{(n)}_{j}$ satisfy:
@@
Solving Eq. (5) by means of the Fourier collocation method, we can obtain the chemical potential as the function of the quasi wave numbers, $\mu_{n}(k_{x},k_{y})$, and the corresponding eigenfunctions $\phi_{j}^{(n)}(x,y,k_{x},k_{y})$ in the first Brillouin zone. Figures 1(c) and 1(d) show the band-gap structures along the high-symmetry lines for the MLs with $\theta=\arctan(3/4)$ and $\theta=\arctan(5/12)$, respectively. Comparing the lowest three energy bands in these two cases, we find that the energy bands for $\theta=\arctan(5/12)$ are relatively flat and have lower values of $\mu_{n}$. However, the energy gaps for $\theta=\arctan(3/4)$ are much wider than those for $\theta=\arctan(5/12)$, especially for the second finite energy gap. The presence of multiple and wide energy gaps enables the system to support various GS types.
The depth and period of the lattice potential significantly affect the band-gap structure, a deeper lattice potential resulting in narrower bands and wider band gaps. We consider the effect of lattice period $a$ on the band-gap structure for $\theta=\arctan(5/12)$, $V_{0}=4$, and $\gamma=0.5$, as shown in Fig. 1(e). As the lattice period $a$ increases, the chemical potential $\mu$ of each energy band decreases. Simultaneously, the increase in $a$ leads, first, to a merger of the high-energy bands, followed by splitting into narrower (nearly flat) bands, accompanied by the disappearance and reemergence of the gaps which separate the bands. The emergence of multiple flat bands here is an ML-produced effect.
We also address the effect of the SOC strength $\gamma$ on the band-gap structure fixing given $\theta=\arctan(5/12)$, $V_{0}=4$, and $a=0.5$, as shown in Fig. 1(f). Similar to the effect of $a$, the increase in $\gamma$ leads to a decrease in the chemical potential of each energy band, but at a faster rate. Due to the varying rates of the decrease for each energy band, the widths of the energy gaps are affected and an additional energy gap forms.
@@ III.2 Wannier-type and Bloch-type gap solitons
We have found bright solitons, i.e., GSs populating the gaps. If both interaction coefficients $c_{0}$ and $c_{2}$ are negative (which corresponds to the attractive nonlinearity), the 2D system suffers the collapses, as might be expected [54, 55], which inhibits the formation of GSs. On the other hand, if both $c_{0}$ and $c_{2}$ are positive (which corresponds to the repulsive nonlinearity), the broadening of wave functions induced by dispersion cannot be balanced by the repulsive interactions. GSs can only exist in a large-scale potential well (binding all atoms) of the ML, which is similar to the harmonic-oscillator trapping potential [56]. Here, we produce GS solutions of Eq. (1) in the most interesting case of $c_{0}c_{2}<0$, i.e., two cases $c_{0}<0$, $c_{2}>0$ (antiferromagnetic), and $c_{0}>0$, $c_{2}<0$ (ferromagnetic) [48]. Here, we fix $\theta=\arctan(3/4)$ to deal with broader energy gaps.
GS solutions $\varphi_{j}(x,y)$ satisfy the stationary-state equations, which are obtained by substituting $\psi_{j}(x,y,t)=\varphi_{j}(x,y)e^{i\mu t}$ into Eq. (1):
@@
GS solutions to Eq. (6) were obtained using the numerical squared-operator iteration method [57], with the computational grid $(x,y)\in[-12\pi,+12\pi]$ and the number of discrete points is $1024\times 1024$. For the ML with $\theta=\arctan(3/4)$, $a=2\pi$, and $V_{0}=4$, the low-energy bands exhibit flat dispersion, and several gaps are present. We have found five types of GSs, setting $\gamma=0.5$, $c_{0}=-1$, and $c_{2}=1.5$ (the antiferromagnetic system). The density profiles of the five GS types are plotted in Figs. 2(a)-2(e), featuring ring-shaped and multi hump patterns.
Figure 2: WT (Wannier-type) GS families bifurcating from the lowest five flat bands for the ML with $\theta=\arctan(3/4)$, $a=2\pi$, and $V_{0}=4$. Other parameters are $\gamma=0.5$, $c_{0}=-1$, and $c_{2}=1.5$ (the antiferromagnetic system). (a)–(e) The density profiles of three components $\varphi_{\pm 1}$ (the first row) and $\varphi_{0}$ (the second row) for five GS families. The corresponding phases are shown in the corresponding subfigures, where the phase of $\varphi_{+1}$ is in the left subfigure. The drawing grid is $(x,y)\in[-3,3]$. (f) Norm NN vs the chemical potential $\mu$ for the five GS families. The solid (dashed) parts of these curves indicate that the corresponding GSs are stable (unstable). The black lines represent five flat bands with $\mu_{1}=1.6998$, $\mu_{2}=1.8065$, $\mu_{3}=3.5533$, $\mu_{4}=3.5808$, and $\mu_{5}=3.7703$. Here, the inset shows the norm curve of the GS family bifurcating from $\mu_{1}$, plotted near the band edge.
Note that the density profiles of the GS components $\psi_{+1}$ and $\psi_{-1}$ are identical, but they have different phases. In Fig. 2(f) we also present the relation between the soliton’s norm, $N=\int_{-\infty}^{+\infty}\sum_{j}|\varphi_{j}|^{2}\mathrm{d}x\mathrm{d}y$, and the chemical potential $\mu$ for the five GS families. The norm curves bifurcate at the chemical potential $\mu_{n}$ ($n=1,2,3,4,5$) of the corresponding flat bands and pass through the lower-energy flat bands toward $\mu\rightarrow$-$\infty$. In other words, the chemical potential of the GS families is
@@
where $\Delta\mu<0$, provided that $c_{0}<0$. It is easy to understood that the offset $\Delta\mu$ is determined by $c_{0}$ because these GS solutions are invariant with respect to the spin-flip symmetry (7), which implies that the nonlinear terms associated with $c_{2}$ are eliminated when substituting the GS solutions into Eq. (6). As $\mu$ increases, the density profiles of these five GS types remain almost unchanged (slightly widened), but norm $N\rightarrow 0$ as $\mu\rightarrow\mu_{n}$. Here, the norm of GSs exhibits a very small threshold on the order of $10^{-8}$, as shown in the inset of Fig. 2. In addition, the square minimal unit cell of the ML gives rise to three other flat bands that are very close to those ($\mu_{2},\mu_{3},\mu_{4}$), and there exist three additional GS types (not shown here) whose density profiles are very similar to those plotted in Figs. 2(b)-2(d), rotated by $\pi/4$. The norm curves of the two degenerate GS types are strongly overlapping. In fact, all these GSs are similar to Wannier functions $w_{j}^{(n)}(x,y)$ (in the central unit cell) corresponding to $\mu_{n}$, especially in the limit of small amplitudes, where the nonlinear terms may be treated as small perturbations, i.e.,
@@
where $S=\pi^{2}/(5a^{2})$ is the area of the first Brillouin zone [58]. The Wannier functions corresponding to the flat bands are eigenfunctions of the linear part of Eq. (1), with the eigenvalue of $\mu_{n}$. This is why these GS families bifurcate at the chemical potential $\mu_{n}$ corresponding to the flat bands. The GSs, which remain localized even at small values of the norm, can be well approximated by Wannier functions; therefore they are termed WT GSs.